riah691:

The triangles are similar. If WZ = 15, ZY = 12, and XY = 11, find WV. Round to the nearest tenth.

BenLindquist:

do you know what the formula is?

riah691:

No, do you?

BenLindquist:

use XY / ZY = WV / WZ

BenLindquist:

for the formula

riah691:

okay! thank you

BenLindquist:

you're welcome

riah691:

i'm supposed to find WV

BenLindquist:

do you have a diagram?

riah691:

yes, hold on

BenLindquist:

you aren't looking for x right?

riah691:

1 attachment
BenLindquist:

what are your answer choices?

riah691:

8.8, 16.4, 15.0, and 13.8

BenLindquist:

XY / ZY = WV / WZ 11/12 = x/15 x = ??

riah691:

i don't know

BenLindquist:

13.8

riah691:

1 attachment
BenLindquist:

can you make a new post for every question?

riah691:

yes!